Dtcelmber 31, 2017 Required Prepare a compalall Analysls Component: Are Derlaks a nced mainly b change in how assets were financed from 2016 to 2017 favourable oq ssets fina uity in 2016? in 2017 n were financed from 2016 to 2017 favourable or u CHECK FIGURES: 1B. 2019 Depreciation expense $39,063 2A. 201q Yorkton Company purchased for ment having an an estimated residual value of $62, a December 31 year-end. Problem 9-3A Calculating depreciation-partial periods LO2.3 nse = $61,524 $375,000 equip 500. Depreciation is calculated to the nearest mo estimated useful life of month.Thecoman Complete the following schedules (round calculations to the ne Required 1. Assuming the purchase was made on January 1, 2017 2. Assuming the purchase was made on July 1, 2017. 2017 2018 2019 A. Double-declining-balance method Equipment. Less: Accumulated depreciation Year-end book value Depreciation expense for the year Equipment Less: Accumulated depreciation Year-end book value Depreciation expense for the year

Please full calculation and explanation.
Answer #1

Solution 1:

Cost of Equipment = $375,000

Residual value = $62,500

Useful life = 8 years

Depreciation rate - SLM = 1/8 = 12.50%

Depreciable cost = Cost - Residual value = $375,000 - $62,500 = $312,500

Annual depreciation - SLM = $312,500 / 8 = $39,063

Depreciation rate - DDB = 12.50%*2 = 25%

Annual depreciation using DDB:

2017 = $375,000*25% = $93,750

2018 = ($375,000 - $93,750) * 25% = $70,313

2019 = ($375,000 - $93,750 - $70,313) * 25% = $52,734

Depreciation Schedule
Particulars 2017 2018 2019
A. Double - Declining balance method:
Equipment $375,000.00 $375,000.00 $375,000.00
Less: Accumulated Depreciation $93,750.00 $164,063.00 $216,797.00
Year End book value $281,250.00 $210,937.00 $158,203.00
Depreciation expense for the year $93,750.00 $70,313.00 $52,734.00
B. Straight line method
Equipment $375,000.00 $375,000.00 $375,000.00
Less: Accumulated Depreciation $39,063.00 $78,126.00 $117,189.00
Year End book value $335,937.00 $296,874.00 $257,811.00
Depreciation expense for the year $39,063.00 $39,063.00 $39,063.00

Solution 2:

Cost of Equipment = $375,000

Residual value = $62,500

Useful life = 8 years

Depreciation rate - SLM = 1/8 = 12.50%

Depreciable cost = Cost - Residual value = $375,000 - $62,500 = $312,500

Annual depreciation - SLM:

2017 = $312,500 / 8 * 6/12 = $19,531

2018 = $312,500/8 = $39,063

2019 = $312,500/8 = $39,063

Depreciation rate - DDB = 12.50%*2 = 25%

Annual depreciation using DDB:

2017 = $375,000*25%*6/12 = $46,875

2018 = ($375,000 - $46,875) * 25% = $82,031

2019 = ($375,000 - $46,875 - $82,031) * 25% = $61,524

Depreciation Schedule
Particulars 2017 2018 2019
A. Double - Declining balance method:
Equipment $375,000.00 $375,000.00 $375,000.00
Less: Accumulated Depreciation $46,875.00 $128,906.00 $190,430.00
Year End book value $328,125.00 $246,094.00 $184,570.00
Depreciation expense for the year $46,875.00 $82,031.00 $61,524.00
B. Straight line method
Equipment $375,000.00 $375,000.00 $375,000.00
Less: Accumulated Depreciation $19,531.00 $58,594.00 $97,657.00
Year End book value $355,469.00 $316,406.00 $277,343.00
Depreciation expense for the year $19,531.00 $39,063.00 $39,063.00